ABC Company employs a periodic inventory system and sells its inventory to customers
for $27 per unit. ABC Company had the following inventory information available for
the month of May:

May 1    Beginning inventory 2,400 units @ $14 cost per unit
May 8    Sold 1,000 units
May 13   Purchased 1,300 units @ $17 cost per unit
May 18   Sold 1,200 units
May 21   Purchased 1,100 units @ $11 cost per unit
May 28   Sold 300 units
May 30   Purchased 700 units @ $10 cost per unit

During May, ABC Company reported operating expenses of $14,500 and had an income tax
rate of 43%.

Calculate the amount of gross profit shown on ABC Company's income statement for May
using the weighted average method.
Answer #1
Units Unit cost Total cost
May 1 2400 14 33600
May 13 1300 17 22100
May 21 1100 11 12100
May 30 700 10 7000
Total 5500 74800
Average cost per unit 13.60 =74800/5500
Total units sold 2500 =1000+1200+300
Sales Revenue 67500 =2500*27
Less: Cost of goods sold 34000 =2500*13.60
Gross Profit 33500